A 560.0 mL of 0.61 M sodium bromide, NaBr solution has a density of 1.046 g mL-1


. Calculate the


molality of the solution.


1
Expert's answer
2021-08-30T02:26:26-0400

"V=560.0 \;mL \\

ρ=1.046 \;g/mL \\

m(solution) = ρV = 1.046 \times 560.0 = 585.76 \;g"

Number of moles of NaBr present in the solution:

Proportion:

0.61 mol – 1000 mL

x mol – 560 mL

"x = \frac{0.61 \times 560}{1000}= 0.3416 \;mol"

Number of moles of NaBr present in the solution is 0.3416

"M(NaBr) = 102.894 \;g/mol \\

m(NaBr) = 0.3416 \times 102.894 = 35.148 \;g"

Mass of water in the solution "= (585.76 -35.148) = 550.612 \;g"

Molality of the NaBr solution "= \frac{0.3416}{550.612} \times 1000 = 0.6204 \;m"